# Self-Normalized Resets for Plasticity in Continual Learning

### v2

@@ 2 Algorithm
Algorithm 1 SNR: Self-Normalized Resets
@@ 2.1 Motivating SNR and Comparison to Other Reset Schemes
Let $\bar{T}$ be the $1-\lambda(p-\lambda)^{-1}$ percentile of a Geometric$(p)$ distribution. Then the optimal hypothesis test takes the form $X_{\tau}=\mathbf{1}\{Z_{\tau}=0\}$ where $\tau={\rm min}(s:Z_{s}=1)\wedge\bar{T}$.
@@ 3.1.1 Algorithms and Architetcures
Our baseline in all problems consist simply of using SGD or Adam as the optimizer with no further intervention. We then consider several interventions to mitigate plasticity loss. First, we consider algorithms that employ an explicit reset of neurons: these include SNR, Continual Backprop (CBP) (Dohare et al., 2021), and ReDO (Sokar et al., 2023). Among algorithms that attempt to use regularization, we consider vanilla L2 regularization, L2 Init (Kumar et al., 2023b), and Shrink and Perturb (Ash & Adams, 2020). Finally, as a potential architectural modification we consider the use of Layer Normalization (Ba et al., 2016).
@@ 3.2 Results and Discussion
We separately discuss the results for the first four problems (PM, RM, RC, CI) followed by Permuted Shakespeare; we observe additional phenomena in the latter experiment which merit separate discussion.
@@
Figure 4: Parameter norms and attention-score entropy on the Permuted Shakespeare experiment.
@@ 3.2.2 Generalization
4 Theoretical Analysis of Learning a Single ReLU
One common hypothesis that seeks to explain plasticity loss is that the network weights obtained from minimizing loss over some task yield poor initializations for a subsequent task; these initializations effectively lead to poor local minima. This section attempts to make this effect transparent through the analysis of the simplest non-trivial task: learning a single ReLU. Vardi et al. (2021) have already established that for random initializations a ReLU can be efficiently learned. Here we ask what happens when the initialization is picked adversarially; one can think of this adversarial initialization as corresponding to optimal weights for an earlier ReLU learning task.
Theorem 4.1.
[Informal restatement of Theorems E.2 and E.1] There exists a general class of target ReLUs, a distribution over example inputs, and a distribution over initial weights for which:
  • Gradient descent with L2 Init or L2 regularization fail to learn the target ReLU with positive probability with respect to the distribution over initial weights. Specifically the average expected MSE exceeds a positive constant as the number of training examples $T$ grows with a probability bounded from below by a positive constant.
  • Gradient descent with oracle resets achieves an average expected MSE of $O(1/T)$ with probability one.
The oracle resets in the second result above effectively know precisely whether or not a neuron will or will not fire over the input distribution and are in this sense oracular. Reset methods must come as close as possible to mimicking such an oracle; Propositions 2.1 and 2.2 make the case for why SNR might be a good such candidate and why competing reset proposals may fall short. See Appendix E for a formal setup, statement of results, and proofs.
@@ 5 Discussion and Limitations
Theoretical Motivation: Our theoretical development showed that regularization methods themselves are insufficient at combatting the plasticity loss problem, at least in the case of learning ReLUs (Theorem 4.1). Further we showed that existing proposals to detect whether it was appropriate to reset a ReLU (based on ad hoc notions of a neuron’s utility) could have high type 1 error rates relative to an optimal resetting mechanism (Proposition 2.2). SNR was designed to minimize the error rate in solving this detection problem (Proposition 2.1).
@@ A.1 Even Larger Scale Experiment for Permuted Shakespeare
The scale of our largest model is comparable to the scale of the largest transformers considered in recent literature on plasticity loss, namely ViT Tiny (5 million parameters) in Lee et al. (2024). The largest dataset on which Lee et al. (2024) train on is the Tiny Imagenet dataset consisting of 100’000 training examples, on which models are trained for 100 epochs; our largest experiment consists of 775’000 training examples. It is important to consider the scale of datasets used in continual learning experiments as the phenomenon of plasticity loss has been shown to be correlated with the amount of data on which a network trains (Kumar et al., 2023a; Dohare et al., 2021).
@@ Appendix C Additional Experimental Details and Hyperparameter Sweep
For each algorithm we vary its hyperparameter(s) by an appropriate constant over 7 choices, effectively varying the hyperparameters over a log scale. With the exception of the Permuted Shakespeare experiment, we limit over hyperparameter search to 5 choices. In Table 7 we provide the hyperparameter sweep for the 4 benchmark problems. CBP’s replacement rate $r$ is to be interpreted as one replacement per layer every $r^{-1}$ training examples, as presented in Dohare et al. (2024). ReDO’s reset frequency $r$ determines the frequency of resets in units of tasks, as implemented and evaluated in Kumar et al. (2023b).
@@
L2 Reg. ($\lambda$) $10^{-6}$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$ $10^{0}$
L2 Init ($\lambda$) $10^{-6}$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-1}$ $10^{0}$

### v3

@@ 3.1.2 Non-Adversarial Target-Weights
Algorithm 2 Gradient Descent for Minimizing $\mathcal{L}(w)$ with Non-Adversarial Target Weights
Vardi et al. (2021) show that gradient descent (Algorithm 2) under mild distributional conditions which capture our setting of $x\sim N(0,I_{d})\times\{1\}$, converges linearly to the unique global minimizer $w^{*}=v$:
Theorem 3.1 (High-Probability Convergence, Corollary 5.5 of Vardi et al. (2021)).
Under the boundedness (Assumption 5.1) and spread (Assumption 5.3) conditions on the first $d$ coordinates, initialize
$w_{0}=(\tilde{w}_{0},\,0),\qquad\tilde{w}_{0}\;\sim\;\mathrm{Uniform}\bigl(\mathbb{S}^{d-1}(\rho)\bigr),\quad\rho=\frac{M}{c^{2}}.$
Then for any step‐size $\alpha\leq\gamma/c^{4}$, there exist constants $M,c,\gamma>0$ (depending on the distribution) such that with probability at least $1-e^{-\Omega(d)}$,
$\|w_{t}-v\|^{2}\;\leq\;\bigl(1-\gamma\,\alpha\bigr)^{t}\,\|w_{0}-v\|^{2}.$
The high-probability, linear-rate convergence established in Theorem 3.1 implies an $O(1/T)$ expected average regret when training a single ReLU neuron with non-adversarial targets. In contrast, we show below that if the target weights can be chosen after the network is initialized, that is adversarially, the average regret no longer vanishes, even under very mild conditions.
3.1.3 Adversarial Target-Weights
In continual learning, the weight initialization for a task can be those learned for a preceding task and can yield a poor initialization or local minimum. We formalize this notion by introducing the setting of adversarially selected target weights $v$. We formally define ($L_{2}$-regularized) gradient descent under adversarial selection of target weights below in Algorithm 3.
Weight Initialization : Initialize $w_{0}$ from some distribution
Target Selection : Target weights $v$ chosen adversarially with knowledge of $w_{0}$
for $t\leftarrow 0,1,2,\dots$ do
   Compute gradient $\nabla\mathcal{L}(w_{t})$
   Update weights: $w_{t+1}\leftarrow w_{t}-\alpha\,\nabla\mathcal{L}(w_{t})$
end for
Algorithm 3 Gradient Descent for Minimizing $\mathcal{L}(w)$ with Adversarial Target Weights
Below in Theorem, 3.2 we show that ($L_{2}$-regularized) gradient descent can attain $\Omega(1)$ average regret under this adversarial model.
Theorem 3.2.
[…]
$\forall T\geq 0,\quad R_{T}\;\geq\;L(0)\;>\;0.$
We interpret this result as $w_{0}$ being the optimal solution for a preceding task and thereby the starting point for the current task at time $t=0$. It is a priori unclear that there is an efficient way to detect when network weights $w_{t}$ are in a poor local minimum, and moreover, how to remedy it. Fortunately, weight-resets in conjunction with a reset-oracle solve this issue.
3.1.4 Adversarial Target Weights with a Reset Oracle
We define the reset oracle $\mathcal{R}_{\delta}(w)$ as
$\mathcal{R}_{\delta}(w)=\begin{cases}1,&\text{if }\mathbb{E}[\sigma(w^{\top}x)]\leq\delta,\\ 0,&\text{if }\mathbb{E}[\sigma(w^{\top}x)]>\delta.\end{cases}$
The reset oracle $\mathcal{R}_{\delta}$ is parameterized by a threshold $\delta$ and returns 1 (True) if and only if the expected activation of the neuron, defined by the weight vector $w$, falls below $\delta$. This reset oracle is an idealization of the reset-criteria of reset-algorithms like CBP, ReDO, and SNR.
Next, we introduce the gradient descent algorithm with the reset oracle $\mathcal{R}_{\delta}$ for minimizing the loss function $L(w)$. The reset oracle monitors the neuron’s expected activation and determines whether a reset is necessary. If the reset condition is met, the weights are reinitialized from the original distribution; otherwise, the standard gradient descent update is applied to minimize $L(w)$.
Weight Initialization : Initialize $w_{0}$ from some distribution
Target Selection : Target weights $v$ chosen adversarially with knowledge of $w_{0}$
for $t\leftarrow 0,1,2,\dots$ do
 if $\mathcal{R}_{\delta}(w_{t})=0$ then
      Sample $w_{t+1}$ from the initial distribution
 else
      Compute gradient $\nabla L(w_{t})$
      Update weights: $w_{t+1}\leftarrow w_{t}-\alpha\,\nabla L(w_{t})$
   end if
end for
Algorithm 4 Gradient Descent with Reset Oracle for Minimizing $L(w)$
Below, we present our main theoretical contributions. When the target weights $v$ are selected adversarially with knowledge of the network initialization $w_{0}$, Theorem 3.3 establishes that gradient descent with the reset oracle (Algorithm 4) achieves vanishing expected average regret.
Theorem 3.3.
Let $w_{0}$ be sampled from $\mathcal{D}_{w_{0}}=\text{Unif}(l\cdot S^{d-1}\times\{0\})$ for some positive constant $l>0$, where $l\cdot S^{d-1}\subseteq\mathbb{R}^{d}$ is the $l$-radius sphere. Let $v\in\mathbb{R}^{d+1}$ (possibly chosen adversarially with knowledge of $w_{0}$) be the target neuron’s weights satisfying
$v_{d+1}\leq\frac{\|v_{1:d}\|}{2\sqrt{2\pi}},\quad\|v_{1:d}\|=2\pi l,\quad\|v\|\geq C_{1},\quad\|v_{1:d}\|\geq C_{2}\|v\|,$
for some constants $C_{1},C_{2}>0$, and suppose
$\delta<\left(\frac{C_{2}^{2}}{8\pi^{2}\Bigl(\sqrt{3}+\frac{2}{C_{1}}\Bigr)}\right)^{3}.$
Then, minimizing $L(\cdot)$ via gradient descent with step size $\alpha\leq\frac{1}{2(d+1)}$, initialized at $w_{0}$, and employing a reset oracle $\mathcal{R}_{\delta}$, i.e. Algorithm 4, yields
$\forall\,T\geq 0,\quad\mathbb{E}[R_{T}]\leq\mathcal{O}\!\Bigl(\frac{d^{2}\ln T}{T}\Bigr),$
where the expectation is taken over the randomness of the the initialization of $w_{0}$ and the resets.
Remarks. For simplicity of exposition, Theorem 3.3 assumes that $w_{1:d}$ is sampled uniformly from the sphere of radius $l=\frac{\|v_{1:d}\|}{2\pi}$. Alternatively, one may sample $w_{1:d}$ from $\mathcal{N}(0,\sigma^{2}I_{d})$ with $\sigma^{2}=\frac{\|v_{1:d}\|^{2}}{4\pi^{2}d}$, in which case standard concentration of measure arguments imply that
$\mathbb{P}\!\Bigl(\|w\|^{2}\in\Bigl[(1-\epsilon)\frac{\|v_{1:d}\|^{2}}{4\pi^{2}},(1+\epsilon)\frac{\|v_{1:d}\|^{2}}{4\pi^{2}}\Bigr]\Bigr)\geq 1-e^{-\Omega(d\epsilon^{2})}.$
Additionally, the assumption $v_{d+1}\leq\frac{\|v_{1:d}\|}{2\sqrt{2\pi}}$ ensures that the bias term $v_{d+1}$ is not arbitrarily negative relative to $\|v_{1:d}\|$. The assumptions of Theorem 3.3, including the initialization distribution of $w_{0}$, the latter bound on $v_{d+1}$, and the constraints on $\|v\|$ and $\|v_{1:d}\|$, are identical, up to the choice of constants, to those used in the gradient descent analysis of Vardi et al. (2021), the most state-of-the-art work on learning a single ReLU-activated neuron in the non-adversarial setting. While the assumptions of Theorem 3.3 align with those in prior work, our analysis of gradient descent with resets takes an independent approach.
@@ 3.2 Motivating SNR and Comparison to Other Reset Schemes
Let $\bar{T}$ be the $1-\lambda(p-\lambda)^{-1}$ percentile of a Geometric(p)(p) distribution. Then the optimal hypothesis test takes the form $X_{\tau}=\mathbf{1}\{Z_{\tau}=0\}$ where $\tau={\rm min}(s:Z_{s}=1)\wedge\bar{T}$.
@@ 4.1.1 Algorithms and Architetcures
Our baseline in all problems consist simply of using SGD or Adam as the optimizer with no further intervention. We then consider several interventions to mitigate plasticity loss. First, we consider algorithms that employ an explicit reset of neurons: these include SNR, Continual Backprop (CBP) (Dohare et al., 2021), and ReDO (Sokar et al., 2023). Among algorithms that attempt to use regularization, we consider vanilla L2 regularization, L2 Init (Kumar et al., 2023b), and Shrink and Perturb (Ash and Adams, 2020). Finally, as a potential architectural modification we consider the use of Layer Normalization (Ba et al., 2016).
@@
ViT: In addition, we benchmark a Vision Transformer (ViT) on the Continual ImageNet problem. Our ViT follows the specifications of Lee et al. (2024) and Lewandowski et al. (2024), using 4×4 input patches, three attention heads, and 12 transformer layers with an embedding dimension $d_{\text{model}}=192$. This architecture comprises approximately 5.3 million non-embedding parameters. We refer to this variant as CI-ViT to distinguish it from the CNN-based experiments.
@@ 4.2 Results and Discussion
We separately discuss the results for the first four problems (PM, RM, RC, CI) followed by Permuted Shakespeare and CI-ViT; we observe additional phenomena in the latter experiments which merit separate discussion.
@@
Figure 4: Weight norms and attention-score entropy on the Permuted Shakespeare experiment.
Finally, we turn to our CI-ViT experiment, which reproduce many of the same behaviors observed in the Permuted Shakespeare benchmark, a consequence of the transformer architecture. As shown in Figures 5, 6, and 7, plasticity loss in CI-ViT is accompanied by steadily growing weight norms in both the KQV projection matrices and the feed-forward MLP layers, by neuron inactivity, and by a collapse in self‐attention entropy. As with the PS experiment, resets alone remain insufficient to prevent plasticity loss when training a ViT on Continual ImageNet. Table 3 shows that CBP, ReDO, and SNR attain average training accuracies of 0.7037, 0.6380, and 0.6555, respectively, over the final epoch of the last 50 tasks, whereas the regularization-based methods L2 and L2-Init achieve 0.8140 and 0.9196. All algorithms except No Intervention limit feedforward weight norm growth due to regularization or neuron-resets. However, No Intervention and the purely reset-based methods experience unbounded weight norm growth of the KQV projection matrices and a collapse of the entropy of the self-attention probabilities. As stated earlier, this is unsurprising as reset-based methods do not target self-attention layers. To remedy this, we consider the variants SNR+L2 and SNR+L2*. We find that SNR-L2*, like in PS, is the most performant algorithm, attaining an average training accuracy of 0.93372 over the final epoch of the last 50 tasks, while its closest competitor, L2-Init attains 0.91955. Similarly to the MLP and CNN based experiments, we observe that for the transformer-based experiments, PS and CI-ViT, there is no consistent closest competitor to SNR+L2* as L2 is its closest competitor in PS and L2-Init is its closest competitor in CI-ViT.
Algorithm Training Accuracy (Std. Dev.)
No Intv. 0.64315 (0.01403)
L2 0.81399 (0.00811)
L2Init 0.91955 (0.00385)
S&P 0.54589 (0.00072)
CBP 0.70365 (0.03138)
ReDO 0.63802 (0.01893)
SNR 0.65547 (0.02421)
SNR+L2 0.87326 (0.00714)
SNR+L2* 0.93372 (0.00920)
Table 3: Average training accuracy on the last epoch of the final 50 tasks, averaged over 5 seeds with standard deviations in brackets, for the CI-ViT experiment.
Figure 5: 50-task moving average of final-epoch training accuracy of the CI-ViT experiment.
Figure 6: Weight norms on the Continual ImageNet experiment.
Figure 7: Attention-score entropy and number of inactive neurons at any timestep, averaged over each task, on the Continual ImageNet experiment. For reference, our ViT model has 9216 neurons in total.
@@ 5 Discussion and Limitations
Theoretical Motivation: Our theoretical development showed that regularization methods themselves are insufficient at combatting the plasticity loss problem, at least in the case of learning ReLUs (Theorems 3.3 and 3.2). Further we showed that existing proposals to detect whether it was appropriate to reset a ReLU (based on ad hoc notions of a neuron’s utility) could have high type 1 error rates relative to an optimal resetting mechanism (Proposition 3.2). SNR was designed to minimize the error rate in solving this detection problem (Proposition 3.1).
@@ A.1 Even Larger Scale Experiment for Permuted Shakespeare
The scale of our largest model is comparable to the scale of the largest transformers considered in recent literature on plasticity loss, namely ViT Tiny (5 million parameters) in Lee et al. [2024]. The largest dataset on which Lee et al. [2024] train on is the Tiny Imagenet dataset consisting of 100’000 training examples, on which models are trained for 100 epochs; our largest experiment consists of 775’000 training examples. It is important to consider the scale of datasets used in continual learning experiments as the phenomenon of plasticity loss has been shown to be correlated with the amount of data on which a network trains [Kumar et al., 2023a, Dohare et al., 2021].
@@ Appendix C Additional Experimental Details and Hyperparameter Sweep
For each algorithm we vary its hyperparameter(s) by an appropriate constant over 7 choices, effectively varying the hyperparameters over a log scale. With the exception of the Permuted Shakespeare experiment, we limit over hyperparameter search to 5 choices. In Table 8 we provide the hyperparameter sweep for the 4 benchmark problems. CBP’s replacement rate $r$ is to be interpreted as one replacement per layer every $r^{-1}$ training examples, as presented in Dohare et al. [2024]. ReDO’s reset frequency $r$ determines the frequency of resets in units of tasks, as implemented and evaluated in Kumar et al. [2023b].
@@
L2 Reg. ($\lambda$) $10^{-6}$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $^{-1} 1001$$^{0}$
L2 Init ($\lambda$) $10^{-6}$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $^{-1} 1001$$^{0}$